# VScan: Rethinking Visual Token Reduction for Efficient Large Vision-Language Models

### v2

@@ 5.4 Efficiency Analysis
LLaVA-1.5-7B 3.817 T 1113 s (1.00×1.$0$) 416 s (1.00×1.$0$) 100.0% 85.9
+ Ours (33%) 1.253 T 937 s (1.19×1.19$\times$1.19 ×) 301 s (1.38×1.38$\times$1.38 ×) 39.9% 86.2
+ Ours (11%) 0.415 T 812 s (1.37×1.37$\times$1.37 ×) 235 s (1.77×1.$7$) 19.9% 85.0
LLaVA-NeXT-7B 20.825 T 2294 s (1.00×1.$0$) 1420 s (1.00×1.$0$) 100.0% 86.5
+ Ours (33%) 6.459 T 1701 s (1.35×1.35$\times$1.35 ×) 994 s (1.43×1.43$\times$1.43 ×) 34.8% 86.1
+ Ours (11%) 2.099 T 1120 s (2.05×2.05$\times$2.05 ×) 488 s (2.91×2.91$\times$2.91 ×) 13.1% 85.1
In Table 6, we evaluate the practical acceleration effects of VScan. By retaining only 11% of the visual tokens, VScan achieves a 1.37$\times$ speedup in overall efficiency and a 1.77$\times$ speedup in prefilling efficiency on LLaVA-1.5-7B, while maintaining robust performance with only a 0.9% decline. Our approach achieves even more significant acceleration on LLaVA-NeXT-7B, where it delivers a 2.05$\times$ speedup in inference and a 2.91$\times$ speedup in the prefill stage. Additionally, our approach can also effectively compress KV cache storage across different backbones. It is also important to note that VScan is compatible with FlashAttention [17], which can further enhance efficiency. For instance, the inference time of our approach with an 11% retention rate on LLaVA-NeXT-7B can be further reduced from 488 to 473 seconds.
@@ 6 Conclusion
In this work, we present a comprehensive empirical study to understand how visual information is processed across both the visual encoding and LLM decoding stages. Building on these insights, we propose VScan—a two-stage, training-free visual token reduction framework—to accelerate LVLM inference while maintaining robust performance. Extensive experiments across 4 LVLM architectures and 16 image and video benchmarks demonstrate that our approach consistently outperforms existing state-of-the-art methods, achieving a superior trade-off between efficiency and accuracy.
@@
A Additional Experimental Results
A.1 Results on Video-LLaVA
Table A1: Performance comparisons on Video-LLaVA-7B [37] across 4 video understanding tasks with a 75% reduction rate. The best results are bolded, and the second-best are underlined. Evaluation is based on our re-implementation.
Method TGIF MSVD MSRVTT ActivityNet
Acc. Score Acc. Score Acc. Score Acc. Score
Video-LLaVA-7B [37] 47.0 3.40 70.5 3.92 58.3 3.51 42.2 3.37
FastV [12] 42.7 3.19 67.4 3.83 53.6 3.40 36.1 3.15
PyramidDrop [60] 44.1 3.26 66.7 3.81 56.1 3.45 37.4 3.15
VScan (Ours) 46.9 3.35 69.8 3.93 57.1 3.48 42.6 3.34
Finally, we validate the effectiveness of our approach on video understanding tasks and compare its performance against other approaches on Video-LLaVA-7B [37], as shown in Table A1. Specifically, we report the accuracy and GPT-evaluated scores for each benchmark to assess the quality of the responses. The results illustrate that with a 25% token budget, our approach maintains nearly 100% of the original Video-LLaVA-7B’s performance, consistently outperforming other methods.
A.2 Empirical Validation of Global and Local Scan
@@
Figure A1: Performance comparisons on AdvGQA and AdvPOPE. We report the results for each approach with 64 visual tokens retained using LLaVA-1.5-7B [39].
@@
To validate the effectiveness of our global and local scan schemes, we construct adversarial subsets for GQA and POPE, namely AdvGQA and AdvPOPE, which contains failure cases similar to those shown in Figure 2 (Left), where the text queries play an important role and relying solely on the global scan to select visual tokens leads to errors. To accurately select these samples, we follow Gandelsman et al. [22] to decompose the image representations and pinpoint the tokens or regions most relevant to the query. Specifically, we utilize both the text and visual encoders of CLIP-ViT-L-336px [50] to identify the visual tokens relevant to the text query as a reference. Two examples of the visual tokens selected by CLIP are shown in Figure 2 (Left). A sample is included in the adversarial set if the response is correct when using the 64 tokens selected by CLIP, but becomes incorrect when using the 64 tokens selected solely by the global scan. Following this, we collected 886 and 515 adversarial samples in AdvGQA and AdvPOPE, respectively.
@@
In Figure A1, we present a performance comparison of incorporating both global and local scans with FastV [12] and PyramidDrop [60], which select visual tokens based on text attention and are expected to handle samples in the adversarial set effectively. We observe that incorporating both global and local scans achieves performance comparable to these text-guided approaches, despite being text-agnostic and selecting important tokens solely based on visual significance. These results validate that combining both scanning strategies and token merging helps preserve the maximum amount of visual information, effectively preventing information loss.
A.3 Qualitative Results
@@
In Figure A2, we present qualitative examples from the RefCOCO benchmark using Qwen-2.5-VL [6]. For each image, we show the model’s predicted bounding boxes in response to different referring expressions, along with visualizations of the retained visual tokens after token pruning. These examples illustrate that our method can accurately localize the target objects described in queries, while significantly reducing the number of visual tokens used for inference. This demonstrates the model’s ability to preserve semantically important information under token-efficient settings.
@@
Figure A2: Qualitative results on RefCOCO benchmark using Qwen-2.5-VL [6]. We present the predicted boxes for 6 different queries on 2 images, along with visualizations of the retained tokens.
@@ B More Experimental Details
B.1 Computational Complexity
We follow PyramidDrop [60] to compute the theoretical floating-point operations (FLOPs) introduced in the LLM decoding layers during the pre-filling stage for processing the visual tokens. Specifically, in each of the $K$ decoding layers, self-attention calculation with a causal mask is applied, followed by multiple feed-forward network (FFN) layers. The total FLOPs can thus be computed as:
@@ B.1 Computational Complexity
B.2 Full Experimental Details
Models. We applied our approach to four popular LVLMs with different architectures to evaluate its general effectiveness. Specifically, we follow previous work in this field to compare performance on LLaVA-1.5-7B [39], which is widely recognized for academic use and maps an image input to 576 tokens, and LLaVA-NeXT-7B [40], which offers enhanced high-resolution visual understanding by representing an image input using up to 2,880 visual tokens. We also include evaluations on Video-LLaVA-7B [37], which extends the framework to handle video input, processing up to 8 frames with 2,048 visual tokens. Furthermore, we are among the first to present experimental results on the recent Qwen-2.5-VL [6] model with various LLM sizes (3B, 7B, 32B), which incorporates dynamic resolution processing to handle images of varying sizes, supporting token counts ranging from 4 to 16,384.
Benchmarks and Metrics. We conduct extensive experiments on 9 standard image understanding benchmarks, including visual question answering benchmarks such as GQA [25], ScienceQA [44], VQAv2 [23], TextVQA [52] and VizWiz [24]; multi-modal reasoning benchmarks such as MMBench [43], MMBench-CN [43], MME [20], and POPE [35]. We also include evaluations on 3 more challenging referring grounding tasks using RefCOCO [30], RefCOCO+ [30], and RefCOCOg [47], and report the accuracy achieved by different approaches. In these grounding tasks, a localization is considered correct if the predicted bounding box has an IoU score of at least 0.5 with the ground truth. Additionally, we evaluate our approach on 4 video question answering benchmarks: TGIF [27], MSVD [10], MSRVTT [61], and ActivityNet [64]. We follow previous work [46; 60] to utilize both accuracy and the ChatGPT score [^2] Evaluated using gpt-3.5-turbo: https://platform.openai.com/docs/models/gpt-3.5-turbo. as key performance metrics for these video-based benchmarks.
Baselines. We compare the performance of our approach with 6 state-of-the-art visual token pruning methods: ToMe [7], FastV [12], SparseVLM [69], HiRED [4], PyramidDrop [60], and VisionZip [62]. (1) ToMe [7], which uses bipartite soft matching to iteratively merge similar tokens within ViT layers; (2) FastV [12], which drops visual tokens in early layers of the LLM, guided by text-oriented attention score; (3) SparseVLM [69], which selects vision-relevant text tokens to evaluate the significance of visual tokens; (4) HiRED [4], which dynamically assigns token budgets to sub-images for high-resolution image inputs; (5) PyramidDrop [60], which divides the LLM into stages and drops a portion of visual tokens at the end of each stage; (6) VisionZip [62], which selects a set of dominant tokens and merges unselected tokens contextually. To ensure a fair comparison, we directly report the results of these baselines from their respective original papers unless stated otherwise.
Implementation Details. We perform local scan at a shallow layer, specifically at $l=6$ for LLaVA-series models and $l=8$ for Qwen-2.5-VL. For LLM-stage pruning, we select the middle layer as $k=16$ for LLaVA-series models and $k=14$ for Qwen-2.5-VL-7B. For Qwen-2.5-VL-3B and -32B, we similarly select the exact middle layer for pruning.
B.3 GPT-Aided Evaluation on Video Understanding Tasks
@@
Instructions: You are an intelligent chatbot designed for evaluating the correctness of generative outputs for question-answer pairs. Your task is to compare the predicted answer with the correct answer and determine if they match meaningfully. Here’s how you can accomplish the task: ##INSTRUCTIONS: - Focus on the meaningful match between the predicted answer and the correct answer. - Consider synonyms or paraphrases as valid matches. - Evaluate the correctness of the prediction compared to the answer. User Input: Please evaluate the following video-based question-answer pair: Question: {question} Correct Answer: {answer} Predicted Answer: {pred} Provide your evaluation only as a yes/no and score where the score is an integer value between 0 and 5, with 5 indicating the highest meaningful match. Please generate the response in the form of a Python dictionary string with keys ‘pred’ and ‘score’, where value of ‘pred’ is a string of ‘yes’ or ‘no’ and value of ‘score’ is in INTEGER, not STRING. DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. For example, your response should look like this: {‘pred’: ‘yes’, ‘score’: 4}.

### v3

@@ 5.4 Efficiency Analysis
LLaVA-1.5-7B 3.817 T 1113 s ($1.00\times$) 416 s ($1.00\times$) 100.0% 85.9
+ Ours (33%) 1.253 T 937 s ($1.19\times$) 301 s ($1.38\times$) 39.9% 86.2
+ Ours (11%) 0.415 T 812 s ($1.37\times$) 235 s ($1.77\times$) 19.9% 85.0
LLaVA-NeXT-7B 20.825 T 2294 s ($1.00\times$) 1420 s ($1.00\times$) 100.0% 86.5
+ Ours (33%) 6.459 T 1701 s ($1.35\times$) 994 s ($1.43\times$) 34.8% 86.1
+ Ours (11%) 2.099 T 1120 s ($2.05\times$) 488 s ($2.91\times$) 13.1% 85.1
[…]
@@ 6 Conclusion
In this work, we present a comprehensive empirical study to understand how visual information is processed across both the visual encoding and LLM decoding stages. Building on these insights, we propose VScan—a two-stage, training-free visual token reduction framework—to accelerate LVLM inference while maintaining robust performance. Specifically, we design complementary global and local scanning strategies to select informative visual tokens that preserve rich visual details during visual encoding, and further refine this token set via middle layer pruning in the LLM decoding stage based on textual relevance. Extensive experiments across 4 LVLM architectures and 16 image and video benchmarks demonstrate that our approach consistently outperforms existing state-of-the-art methods, achieving a superior trade-off between efficiency and accuracy.
Limitations. One potential limitation of this work is the inherent trade-off between efficiency and accuracy: while the proposed VScan significantly reduces inference cost of LVLMs, aggressive token pruning may still distort visual information and lead to degraded performance, particularly on challenging tasks that demand fine-grained understanding or compositional reasoning.
@@
A Additional Experimental Results and Discussions
A.1 Empirical Validation of Global and Local Scan
@@
Figure A1: Performance comparisons on AdvGQA and AdvPOPE. We report the results for each approach with 64 visual tokens retained using LLaVA-1.5-7B Liu et al. [2024a].
@@
To validate the effectiveness of our global and local scan schemes, we construct adversarial subsets for GQA and POPE, namely AdvGQA and AdvPOPE, which contains failure cases similar to those shown in Figure 2 (Left), where the text queries play an important role and relying solely on the global scan to select visual tokens leads to errors. To accurately select these samples, we follow Gandelsman et al. Gandelsman et al. [2024] to decompose the image representations and pinpoint the tokens or regions most relevant to the query. Specifically, we utilize both the text and visual encoders of CLIP-ViT-L-336px Radford et al. [2021] to identify the visual tokens relevant to the text query as a reference. Two examples of the visual tokens selected by CLIP are shown in Figure 2 (Left). A sample is included in the adversarial set if the response is correct when using the 64 tokens selected by CLIP, but becomes incorrect when using the 64 tokens selected solely by the global scan. Following this, we collected 886 and 515 adversarial samples in AdvGQA and AdvPOPE, respectively.
@@
In Figure A1, we present a performance comparison of incorporating both global and local scans with FastV Chen et al. [2024a] and PyramidDrop Xing et al. [2025], which select visual tokens based on text attention and are expected to handle samples in the adversarial set effectively. We observe that incorporating both global and local scans achieves performance comparable to these text-guided approaches, despite being text-agnostic and selecting important tokens solely based on visual significance. These results validate that combining both scanning strategies and token merging helps preserve the maximum amount of visual information, effectively preventing information loss.
A.2 Qualitative Results
@@
In Figure A2, we present qualitative examples from the RefCOCO benchmark using Qwen-2.5-VL Bai et al. [2025]. For each image, we show the model’s predicted bounding boxes in response to different referring expressions, along with visualizations of the retained visual tokens after token pruning. These examples illustrate that our method can accurately localize the target objects described in queries, while significantly reducing the number of visual tokens used for inference. This demonstrates the model’s ability to preserve semantically important information under token-efficient settings.
@@
Figure A2: Qualitative results on RefCOCO benchmark using Qwen-2.5-VL Bai et al. [2025]. We present the predicted boxes for 6 different queries on 2 images, along with visualizations of the retained tokens.
A.3 Results on More Real-World Benchmarks
To comprehensively assess the effectiveness of VScan, we evaluate its performance on a range of real-world benchmarks, including DocVQA, InfoVQA, MME-RealWorld, MM-Vet, and MMMU, using the LLaVA-1.5-7B baseline. As shown in Table A1, VScan achieves competitive results while significantly reducing the number of visual tokens. With only 192 tokens, VScan maintains performance nearly identical to the full-token LLaVA-1.5-7B baseline, even surpassing it on MM-Vet and MMMU. When further reducing tokens to 128 and 64, VScan still preserves strong accuracy, demonstrating graceful degradation across tasks. These results highlight the robustness and efficiency of our method, showing that substantial token reduction can be achieved without sacrificing performance on diverse real-world benchmarks.
We further compare VScan with FastV and PDrop under different token retention rates on DocVQA and InfoVQA, using both LLaVA-NeXT-7B and Qwen2.5-VL-7B backbones. As shown in Table 2(b), VScan consistently outperforms FastV and PDrop at the same retention levels. For LLaVA-NeXT-7B, VScan achieves 72.6 on DocVQA and 36.3 on InfoVQA with only 33% of tokens, closely matching the full-model baseline (74.4/37.1) while retaining higher accuracy than FastV and PDrop. Similar trends hold at more aggressive pruning levels, where VScan shows lower degradation compared to stronger drops for the baselines. Results on Qwen2.5-VL-7B further confirm this advantage: VScan reaches 94.0/79.7 at 33% retention and maintains 83.9/73.4 even at 11%, substantially outperforming FastV and PDrop. These results demonstrate that VScan generalizes across architectures and provides a more robust trade-off between efficiency and accuracy.
Table A1: Performance of our VScan across diverse real-world benchmarks. We report results on DocVQA, InfoVQA, MME-RealWorld, MM-Vet, and MMMU. Compared with the LLaVA-1.5-7B baseline, VScan achieves competitive performance while substantially reducing the number of visual tokens retained.
Method Retention Rate DocVQA InfoVQA MME-RealWorld $M$-Vet MMMU
LLaVA-1.5-7B 28.1 25.8 24.9 31.1 35.3
VScan 33.3% 27.5 25.7 24.0 31.8 35.7
VScan 22.2% 25.7 25.4 22.7 30.5 36.1
VScan 11.1% 23.9 23.7 22.2 29.7 35.8
Table A2: Comparison of VScan with FastV and PDrop on DocVQA and InfoVQA. Results are reported under different token retention rates (33%, 22%, 11%). Evaluation is based on our re-implementation.
Method DocVQA InfoVQA
LLaVA-NeXT-7B 74.4 37.1
FastV (33%) 69.8 33.5
PDrop (33%) 71.8 35.0
VScan (33%) 72.6 36.3
FastV (22%) 64.7 32.8
PDrop (22%) 67.4 33.4
VScan (22%) 71.6 34.2
FastV (11%) 61.2 28.1
PDrop (11%) 64.0 31.6
VScan (11%) 68.8 34.5
(a) LLaVA-NeXT-7B.
Method DocVQA InfoVQA
Qwen2.5-VL-7B 95.7 82.6
FastV (33%) 88.7 75.3
PDrop (33%) 90.1 76.2
VScan (33%) 94.0 79.7
FastV (22%) 81.6 70.8
PDrop (22%) 83.8 73.7
VScan (22%) 87.8 77.1
FastV (11%) 74.9 65.2
PDrop (11%) 77.4 67.0
VScan (11%) 83.9 73.4
(b) Qwen2.5-VL-7B.
A.4 VScan for Accelerating Training
We further validate the effectiveness of VScan in reducing training cost while preserving the model performance. As shown in Table A3, VScan reduces training time by 41.7% from 96 GPU hours to 56 GPU hours, achieving 96.7% of the original performance. For comparison, we also report inference-time VScan results, which reach 95.2% of baseline performance. Notably, applying VScan during training yields a 1.5% improvement over using VScan solely at inference.
Table A3: Training efficiency and performance of VScan compared with the LLaVA-1.5-7B baseline. VScan significantly reduces GPU hours while maintaining strong performance across benchmarks.
Method GPU Hours GQA MMB MMB-CN MME POPE Avg.
LLaVA-1.5-7B 96 61.9 64.7 58.1 1862 85.9 100.0%
VScan (fine-tune) 56 59.2 63.0 57.9 1721 84.6 96.7%
VScan (inference) 58.3 62.1 55.7 1698 85.0 95.2%
A.5 Remarks on Multi-Turn Conversations
Adapting the middle-layer pruning component of VScan to support multi-turn conversations is straightforward: When presented with new questions, VScan can reassess token importance and reselect textually relevant visual tokens from the existing token pool through global-local scans. Although this re-selection introduces additional computation compared to text-agnostic methods (which are inherently compatible with multi-turn dialogue), the overhead is negligible relative to the overall prefill time.
A.6 More Efficiency Comparisons
Table A4 summarizes the trade-off between efficiency and performance for various visual-token pruning methods across different retention ratios (33.3%, 22.2%, and 11.1%). VScan consistently achieves superior accuracy while maintaining competitive FLOPs and inference time, demonstrating its effectiveness in preserving visual information under aggressive pruning.
Table A4: Efficiency and performance comparison under different token retention ratios. We report the total FLOPs (T), inference time (s), and accuracy achieved by the model at a given retention rate.
Method 33.3% 22.2% 11.1%
LLaVA-1.5-7B - 3.817T / 416s / 85.9 -
FastV 1.253T / 298s / 60.7 0.836T / 266s / 57.2 0.421T / 231s / 44.5
PDrop 1.253T / 307s / 82.3 0.834T / 278s / 82.3 0.415T / 240s / 55.9
VisionZip 1.253T / 293s / 85.3 0.834T / 263s / 83.2 0.415T / 229s / 77.0
VScan 1.253T / 301s / 86.2 0.834T / 274s / 86.1 0.415T / 235s / 85.0
A.7 Full Numerical Results on Qwen-2.5-VL-7B
We report the full numerical results on Qwen-2.5-VL-7B in Tables A5, A6, and A7.
Table A5: MMBench accuracy under different visual token retention ratios. Higher is better.
Model Method 11% 22% 33% 50% 100%
32B FastV 70.13 78.09 81.01 85.40 87.37
PDrop 73.22 79.04 82.74 86.03 87.37
Ours 78.09 82.47 85.31 86.17 87.37
7B FastV 69.07 74.74 77.73 80.27 84.19
PDrop 71.08 75.43 78.04 81.96 84.19
Ours 76.89 80.50 80.67 82.56 84.19
3B FastV 56.79 67.27 72.42 74.98 78.87
PDrop 55.93 63.22 70.30 72.48 78.87
Ours 69.67 70.62 72.81 75.64 78.87
Table A6: MMBench-CN accuracy under different visual token retention ratios. Higher is better.
Model Method 11% 22% 33% 50% 100%
32B FastV 67.61 76.12 79.81 83.42 86.00
PDrop 71.31 77.65 80.72 84.28 86.00
Ours 77.23 80.76 84.45 84.78 86.00
7B FastV 66.92 70.53 76.98 77.41 81.79
PDrop 68.59 69.85 76.12 78.15 81.79
Ours 74.48 76.37 78.69 79.04 81.79
3B FastV 54.73 66.32 71.13 74.05 77.41
PDrop 57.20 65.92 71.58 74.79 77.41
Ours 63.66 68.47 72.45 74.51 77.41
Table A7: MME total scoresunder different visual token retention ratios. Higher is better.
Model Method 11% 22% 33% 50% 100%
32B FastV 1536 1924 2124 2316 2432
PDrop 1536 1797 2078 2279 2432
Ours 2015 2189 2288 2378 2432
7B FastV 1970 1990 2240 2257 2291
PDrop 1982 2064 2218 2236 2291
Ours 1990 2135 2275 2270 2291
3B FastV 1620 1800 1953 2053 2139
PDrop 1444 1692 1709 1986 2139
Ours 1763 1838 1977 2037 2139
@@ B More Experimental Details
B.1 Remarks on Ensuring Fair Comparisons
In our main result table, we ensured fair comparisons between our VScan and text-agnostic approaches such as VisionZip. Instead of aligning on the final token retention ratio, we matched the average token retention across all LLM layers, which roughly correlates with the inference complexity/speed. For example, at 11.1% retention rate, VScan retains 96 tokens before the LLM using a global-local scan strategy and then prunes to 32 tokens at the middle LLM layer (layer 16 of 32). This yields an average of 64 tokens retained across all LLM layers—matching the average retention of the text-agnostic methods, ensuring that inference speed is consistent.
Additionally, in the comparisons, the compared text-aware approaches, such as FastV and SparseVLM, actually have an advantage over our method, as their reported results are based on the final token count. Since we were unable to fully reproduce all of these methods, we directly reported their results as presented in their respective papers.
To account for this, in the Qwen-2.5-VL and Video-LLaVA experiments presented, we reproduce FastV and PyramidDrop while aligning the average token retention across all LLM layers to ensure a fair comparison (marked in the tables with a †). For instance, for FastV, to achieve an average retention rate of 11%, we prune to 5.2% of the original token count starting from LLM layer 2, resulting in an average retention of (100% × 2 + 5.2% × 30) / 32 = 11.1%.
Therefore, under the same average retention rate, the efficiency metrics are essentially similar. This means that, at the same retention rate, higher performance directly translates into a more favorable performance–efficiency trade-off. In this regard, we have conducted thorough experiments demonstrating that our approach achieves a state-of-the-art performance–efficiency trade-off across benchmarks and models.
B.2 Computational Complexity
We follow PyramidDrop Xing et al. [2025] to compute the theoretical floating-point operations (FLOPs) introduced in the LLM decoding layers during the pre-filling stage for processing the visual tokens. Specifically, in each of the $K$ decoding layers, self-attention calculation with a causal mask is applied, followed by multiple feed-forward network (FFN) layers. The total FLOPs can thus be computed as:
@@ B.2 Computational Complexity
B.3 Experimental Design for Study 3
The goal of Study 3 is to determine at which LLM layers the next-token predictions stabilize, i.e., when the model’s forward computation becomes semantically saturated. Once the prediction has converged, further layers mainly refine numerical precision rather than introduce new semantic information. Understanding this stabilization point is crucial because pruning before convergence can disrupt semantic formation, whereas pruning after convergence is largely safe and provides meaningful computational speedups.
Here is our experimental setup: For each decoding layer $l$ in the LLaVA-1.5-7B LLM backbone, we extract its hidden state $h_{l}$ and pass it through the original final linear projection layer to obtain the logits, i.e., $s_{l}=W_{\text{proj}}h_{l}$. We then apply a softmax to compute the prediction confidence and report the top-confidence token for each layer.
B.4 GPT-Aided Evaluation on Video Understanding Tasks
@@
You are an intelligent chatbot designed for evaluating the correctness of generative outputs for question-answer pairs. Your task is to compare the predicted answer with the correct answer and determine if they match meaningfully. Here’s how you can accomplish the task:
##INSTRUCTIONS:
- Focus on the meaningful match between the predicted answer and the correct answer.
- Consider synonyms or paraphrases as valid matches.
- Evaluate the correctness of the prediction compared to the answer.
User Input:
Please evaluate the following video-based question-answer pair:
Question: {question}
Correct Answer: {answer}
Predicted Answer: {pred}
Provide your evaluation only as a yes/no and score where the score is an integer value between 0 and 5, with 5 indicating the highest meaningful match. Please generate the response in the form of a Python dictionary string with keys ‘pred’ and ‘score’, where value of ‘pred’ is a string of ‘yes’ or ‘no’ and value of ‘score’ is in INTEGER, not STRING. DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. For example, your response should look like this: {‘pred’: ‘yes’, ‘score’: 4}.